OpenStudy (anonymous):

How do I find the missing side of an isosceles triangle if A=42 and B=42. How do I find C? A. 42° B. 69° C. 74° D. 132°

OpenStudy (anonymous):

did you leave something out?

OpenStudy (mathstudent55):

In a triangle, if two sides are congruent, then the opposite angles are congruent. This can be restated as: The base angles of an isosceles triangle are congruent.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):



OpenStudy (anonymous):

So m<C = 42?

OpenStudy (mathstudent55):

The sum of the measures of the 3 angles is 180 deg. x + x + 42 = 180

OpenStudy (aric200):

A triangle must=180Deg

OpenStudy (mathstudent55):

No. m<C is not 42. Solve the equation below. 2x + 42 = 180 Solve for x.

OpenStudy (aric200):

If <C=42 then <B Would have to= 42 as <C and B are = To each other Meaning <A <B and <C together would = 126 deg which couldnt be possible, Dont just throw out random numbers Look at the picture surly <A Could not= <C

OpenStudy (aric200):

I am going to put in in simple terms. O kay <C and <B must be the same or = each other. We know <A is 42 Degrees and can not= the same as <C and B. So to Find Do 180-42 Because a triangle= 180 and we already know m<A You will get 138 Now Divide by 2 because there are 2 angles left/ <C and <B You will get 69 So < C and <B both = 69 <C =69

OpenStudy (anonymous):

Thank you @aric200 ! :)

OpenStudy (mathstudent55):

2x + 42 = 180 2x = 138 x = 69